BETAPHARM PROCUREMENT B. Malic Acid.The electronic auction for malic acid was producing very intriguing results. Betapharm had first requested a sealed bid for a small amount of product from the potential suppliers to learn about the price structure of the various players. Because the spread in prices was large (see Exhibit 4 in (A) case), Wilkenson decided to run an electronic auction before doing further due diligence on the suppliers. He wanted to see the proposed prices of the potential suppliers when they were bidding on a higher volume of business (20,000 kilograms rather than 650 kilograms for the sealed bid). Hopefully, the larger volume would lead to economies of scale and, thus, lower prices for Betapharm. Now he watched as the Taiwanese supplier, Chemlab Pharmaceuticals, reduced its auction bid price by 87%; this supplier was no longer the most expensive. Likewise, Liou Laboratories and Zhang Huang dropped their price (Exhibit 2). Although they did not know the names of competing firms, these suppliers were approaching the bid of the Chinese supplier, Chun Wu. Why wasn’t the other Taiwanese supplier, Yu Sheng, also reducing its price? Yu Sheng appeared to be keeping its price within a reasonable margin of its sealed bid price. Despite the rumors that Chinese suppliers had lower quality standards and could not meet regulatory scrutiny, apparently higher-quality Taiwanese firms were matching their bids. Brannigan and Wilkenson were puzzled by this observation.

CHUN WU FINAL BID = $70 per kg, TAO = $145 per kg

Question 4: Supply Chain Management Betapharm contacted Chun Wu, the lowest-bid supplier for detailed information on its order fulfillment policies. For a high volume of business between 20,000-30,000 kg, you obtain the following: Chun Wu, Tao: Lead Time 7 days, 3 days; Order quantity 350 kg, 210 kg. While Betapharm can place an order from the supplier at any time, due to the capacity constraints of the suppliers, the amount ordered each time must be fixed (350 kg from Chun Wu and 210 kg from Tao). Also, because Chun Wu has limited cash flow, payments from the buyer are required upon order placement, whereas Tao is willing to ship material C.O.D. (cash on delivery). In Betapharm’s Exelon production facility, the daily demand for Malic acid is normally distributed with an average of 70 kg and a standard deviation of 15 kg. Since Exelon is a life saving drug, Betapharm strives to achieve a 99.9% service level for their Malic acid stock. Use the price per kg currently charged by Tao to calculate the current annual inventory holding costs associated with contracting with Tao, and then use Chun Wu’s final bid price to see what it would be with Chun Wu. Hints: • Assume Betapharm operates 300 days per year. • Pharmaceutical-grade Malic acid is highly perishable, and as a result annual inventory holding costs are estimated to be 80% of the value of inventory. • The unit cost Malic acid procured from Chun Wu and Tao depends on the final bid that they submit (see Case Part (B) Exhibit 2).
Answer #1

Answer to Question 4.

For Malic Acid procurement from Tao ---

Max Inventory level = Daily maximum usage X Lead time days = (70 + S.D. 15) kg/day X 3 days

= 85 X 3 kg = 255 Kg

Min Inventory level = 0

Average Inventory level = 255/2 kg = 127.5 kg

Cost of average inventory = 127.5 X 145 = $18,487.50

The Annual Inventory Holding cost = 80% of inventory value (as given in question) =

80% of $18,487.50 = $14,790

For Malic Acid procurement from Chun Wu ---

Max Inventory level = Daily maximum usage X Lead time days = (70 + S.D. 15) kg/day X 7 days

= 85 X 7 kg = 595 Kg

Min Inventory level = 0

Average Inventory level = 595/2 kg = 297.5 kg

Cost of average inventory = 297.5 X 70 = $20,825

The Annual Inventory Holding cost = 80% of inventory value (as given in question) =

80% of $20,825 = $16,660